OpenStudy (anonymous):

Point slope form y+4=-4(x-2) What is the equation in standard form of a parallel line that passes through (0,-2)?

OpenStudy (anonymous):

\[ y + 4 = -4x + 8\]\[4x + y -4=0\]

OpenStudy (anonymous):

On the parallel line, find the ordered pair where x=-2

OpenStudy (anonymous):

SInce....the Line...is Parallel they will have Same Slope m = -a/b = -4 nw using one point Form y-y1 = m(x-x1) y + 2 = -4 ( x -0) y + 2 = -4x + 0 4x + y + 2 =0

OpenStudy (anonymous):

hope...This helps